Two masses collide elastically (hit & bounce) where m1 = 0.5 kg, m2 = 1.5 kg,...

  • Two masses collide elastically (hit & bounce) where m1 = 0.5 kg, m2 = 1.5 kg, v1 = 1 m/s, v2 = 0 m/s
  • Calculate the speeds of the balls after the collisions by using the formulas for elastic collisions:

v1' = [v1 * (m1-m2) / (m1+m2)] + [v2 * (2m2) / (m1+m2)]

v2' = [v1 * (2m1) / (m1+m2)] - [v2 * (m1-m2) / (m1+m2)]
Answer #1

Here,mass of one object=m1=0.5 kg and its velocity v1=1 m/s. Mass of second object=m2=1.5 kg, and its velocity v2=0 m/s

Let the final velocity of one object be v1' and final velocity of second object be v2'

In the problem, it is asked to calculate final velocity using the formula given in the problem.

So, According to formula given in problem: v1'=v1*(m1-m2)/(m1+m2)+v2*(2m2)/(m1+m2)

=1(0.5-1.5)/(0.5+1.5)+0*(2*1.5)/(0.5+1.5)= - 0.5 m/s (negative sign represents direction of final velocity is opposite to that of initial velocity)

Also,According to formula given in problem: v2'=v1(2m1)/(m1+m2) - v2(m1-m2)/(m1+m2)

= 1*(2*0.5)/(0.5+1.5)-0*(0.5-1.5)/(0.5+1.5)=0.5 m/s(positive sign represents direction of final velocity is same as that of initial velocity)